OpenStudy (anonymous):

Solve using natural logarithms: \[3e^{2x}+ 5 = 27 \] Minus five from each side: \[3e^{2x} = 22 \] Divide each side by three: \[e^{2x} = 7.3333\] Now what? I forgot :O

OpenStudy (anonymous):

I think we use ln to cancel the 3? 2x = 1.992430164690206162083930 Divide each side by two: x = 0.9962150823451030810419650

OpenStudy (anonymous):

Possibly? I'm not sure :/

OpenStudy (loser66):

e^2x = 22/3 ln both sides, ln e ^ 2x = ln (22/3) --> 2x = ln (22/3) x = (ln(22/3))/2

OpenStudy (loser66):

exactly what you got, hihihi = 0.996

OpenStudy (anonymous):

Thanks :)